Kingsport Containers Company makes a single product that is subject to wide seasonal variations in demand. The company uses a job-order costing system and computes plantwide predetermined overhead rates on a quarterly basis using the number of units to be produced as the allocation base. Its estimated costs, by quarter, for the coming year are given below:

   

Quarter
   First Second Third Fourth
Direct materials $ 160,000 $ 80,000 $ 40,000 $ 120,000
Direct labor 80,000 40,000 20,000 60,000
Manufacturing overhead 240,000 216,000 204,000 ?
Total manufacturing costs (a) $ 480,000 $ 336,000 $ 264,000 $ ?
Number of units to be produced (b) 80,000 40,000 20,000 60,000
Estimated unit product cost (a) ÷ (b) $ 6.00 $ 8.40 $ 13.20 $ ?

Management finds the variation in quarterly unit product costs to be confusing and difficult to work with. It has been suggested that the problem lies with manufacturing overhead because it is the largest element of total manufacturing cost. Accordingly, you have been asked to find a more appropriate way of assigning manufacturing overhead cost to units of product.

Required:

1. Assuming the estimated variable manufacturing overhead cost per unit is $0.60, what must be the estimated total fixed manufacturing overhead cost per quarter?

2. Assuming the assumptions about cost behavior from the first three quarters hold constant, what is the estimated unit product cost for the fourth quarter?

3. What is causing the estimated unit product cost to fluctuate from one quarter to the next?

4. Assuming the company computes one predetermined overhead rate for the year rather than computing quarterly overhead rates, calculate the unit product cost for all units produced during the year.
Answer #1

Solution 1:

Fixed manufacturing overhead = Total manufacturing overhead - Variable manufacturing overhead

= $240,000 - (80000*$0.60) = $192,000

Solution 2:

Estimated unit product cost for 4th quarter = Total manufacturing costs / Nos of units produced

= ($120,000 + $60,000 + 60000*$0.60 + $192,000) / 60000 = $6.80 per unit

Solution 3:

The fixed portion of manufacturing overhead cost is causing the unit product cost to fluctuate. The unit product cost increases as the level of production decreases because the fixed overhead is spread over fewer units.

Solution 4:

Total manufacturing costs for the year = $480,000 + $336,000 + $264,000 + $408,000 = $1,488,000

Total units produced = 80000 + 40000 + 20000 + 60000 = 200000

Unit product cost = $1,488,000 / 200000 = $7.44 per unit